OpenStudy (anonymous):

fing f(-3) for f(x)= 2/5x-2

OpenStudy (anonymous):

f(-3)=2/5(-3)-2 =2/-15-2 =2/-17

OpenStudy (amorfide):

you have a function containing x f(x)= 2/5x-2 but you want a function containing -3 you have an f(x) so if you had f(-3) where ever there was an x in f(x) you would now have a -3 in that place instead so you replace all x's with -3

OpenStudy (anonymous):

Woops, should there or should there not be parenthesis around the denominator?